OpenStudy (anonymous):

Two cars are 210 miles apart. They leave at the same time, traveling towards each other and meet in 3 hours. If the rate of one car is 10 miles per hour more than the other, what is the rate of the slower car?

mathslover (mathslover):

\( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{Ambassadors (people with A next to their name) are here to help you.}\\ \textbf{If you have any questions about the site, feel free to ask us!}\\ \textbf{Users who appear in Purple are Moderators and help run the site.}\\ \textbf{If you have a serious issue with the site or with the behavior of other}\\ \textbf{users}\\ \textbf{please feel free to raise it with them.}\\ \textbf{Want to study a particular subject? Click 'Find More Subjects'}\\ \textbf{search for it there.}\\ \textbf{If you get a good response from a user , click on best response}\\ \end{array}\) Please also read our Code of Conduct : http://openstudy.com/code-of-conduct

mathslover (mathslover):

3(n+n+10)=210 6n+30=210 6n=180 n=30 The slower car is going 30 mph.. may be this help u :)

mathslover (mathslover):

Note : n is the velocity of the slower car .. i just converted the word problem to equation and then solved for n ..